OpenStudy (anonymous):

Suppose P(x,y) i + Q(x,y) j is a two dimensional vector field. Show that F is irrotational if dP/dY = dQ/dX

ganeshie8 (ganeshie8):

basically you want to show curl = 0

OpenStudy (anonymous):

alright, but can you help me solve it?

ganeshie8 (ganeshie8):

\(Curl(F) = \frac{\partial Q}{\partial x} - \frac{\partial P}{\partial y} = 0 \)

OpenStudy (anonymous):

can you show me the steps to where you got that?

ganeshie8 (ganeshie8):

thats the definition of Curl \(Curl(F) = \frac{\partial Q}{\partial x} - \frac{\partial P}{\partial y} \)

ganeshie8 (ganeshie8):

Since you're given that both partials are equal, the Curl equals 0. since Curl = 0, there wont be any rotation component

OpenStudy (anonymous):

alright

ganeshie8 (ganeshie8):

clearly that doesnt look very exciting, I think you're looking for proving it without using definition of Curl ?

ganeshie8 (ganeshie8):

@eliassaab @sourwing

OpenStudy (anonymous):

not sure really but its a 2 mark question, so i think its correct

ganeshie8 (ganeshie8):

ahh ok :) then just show that Curl of vector field equals 0, so the force cannot cause any rotation effects

OpenStudy (anonymous):

alright bro thanks for your help, it turned out easier than i expected

OpenStudy (anonymous):

its almost suspicious

ganeshie8 (ganeshie8):

hhaha okay, np :)

OpenStudy (anonymous):

:)